OpenStudy (anonymous):

An object accelerates from rest to 85 m/s over a distance of 36m . What acceleration did it experience ?

OpenStudy (anonymous):

\[V^2=2*a*S\] (85)^2=2*a*36 a= (85)^2/72=100.34m/s^2

OpenStudy (anonymous):

use equation \[v^2-u^2=2as\] put u=0,s=36,v=85